# Quantifying disorder through sharp spectroscopic fingerprints in an incompressible magnetic state

### v1

@@
Quantifying disorder through sharp spectroscopic fingerprints in an incompressible magnetic state
(June 9, 2025)
@@
(June 9, 2025)
Disorder significantly impacts the electronic properties of conducting quantum materials by inducing electron localization and thus altering the local density of states and electric transport. In insulating quantum magnetic materials the effects of disorder are less understood and can drastically impact fluctuating spin states like quantum spin liquids. In the absence of transport tools, disorder is typically characterized using chemical methods or by semi-classical modeling of spin dynamics. This requires high magnetic fields that may not always be accessible. Here, we show that magnetization plateaus—incompressible states found in many quantum magnets—provide an exquisite platform to uncover otherwise undetectable amounts of disorder, regardless of the origin of the plateau. Using optical magneto-spectroscopy on the Ising-Heisenberg triangular-lattice antiferromagnet K2Co(SeO3)2 exhibiting a 1/3 magnetization plateau, we identify sharp spectroscopic lines, the fine structure of which serves as a hallmark signature of disorder. Through analytical and numerical modeling, we show that these fingerprints not only enable us to quantify minute amounts of disorder but also reveal its nature—as dilute vacancies. Remarkably, this model explains all details of the thermomagnetic response of our system, including the existence of multiple plateaus. Our findings provide a new approach to identifying disorder in quantum magnets.
@@
The role of disorder in quantum magnetic materials is less understood yet it is crucial for a wide range of systems, particularly quantum spin-liquids (QSLs). These systems are believed to host topological order and itinerant fractional excitations without exhibiting clear, conventional order-parameter-like signatures. The response, survival, and nature of potential QSLs in the presence of chemical disorder are not fully understood. Increasingly, researchers are recognizing that disorder can mimic the expected behavior of a QSL [3, 4, 5, 6, 7, 8]. But a significant challenge lies in quantifying the level of disorder using chemical techniques and translating this information into a model magnetic Hamiltonian, $\mathcal{H}$, whose emergent quantum many-body phases are not known in advance. A successful approach involves fitting inelastic neutron scattering or optical spectroscopy measurements under sufficiently high magnetic fields to polarize the system. In this semi-classical regime, it becomes possible to directly measure $\mathcal{H}$ [9], including the level of chemical disorder [10]. However, practical limitations arise from the need for large magnetic fields. Furthermore, spin-anisotropic systems of current interest may only reach saturation asymptotically [11].
@@
The complex magnetization process of antiferromagnets may offer an alternative strategy. By utilizing quantized magnetization plateaus—when they occur—we show that it is not only possible to extract $\mathcal{H}$ at a much lower magnetic field, but these plateaus also exhibit unique, sharp spectroscopic signatures allowing for the detection of disorder with metrological precision. Similar to the QHE, this sensitivity emerges from the distinctive structure of incompressible magnetic states with quantized magnetization, an effect which has not been widely acknowledged until now.
@@
Figure 1: Magnetization plateau and far-infrared magneto-spectroscopy measurements of K2Co(SeO3)2. a, Schematic representation of the normalized uniform magnetization $m=\langle S_{\rm tot}^{z}\rangle/NS$ of an ordered antiferromagnet described at the classical level (solid line) or including quantum corrections (dashed line). The spin-$S$ moments form a canted structure and become fully polarized at the saturation magnetic field $H_{\rm sat}$. b, Normalized uniform magnetization of an antiferromagnet with a field-induced quantized plateau state, which is stabilized between $H_{\rm c1}$ and $H_{\rm c2}$. The UUD state is a possible classical state realizing a $m=1/3$ magnetization plateau. c, Crystal structure of KCSO. Orange spheres represent oxygen, purple represents potassium, blue represents cobalt, and yellow represents selenium ions. The triangular lattice of magnetic Co2+ ions is represented by blue bonds, with the Ising-like nearest-neighbor exchange interactions represented by a brown ellipsoid. d, Frequency and magnetic field dependence of the normalized magneto-transmission spectra $I(\omega;H)$ of KCSO measured in Faraday geometry at $T\!=\!5$ K with the magnetic field direction $z$ along the $c$-axis, see Methods and Supplementary Information (SI) Sec. IA–B for details of our measurement and data analysis protocols. e, Energy modes $E(H)$ extracted from Lorentzian fits to the spectra at each magnetic field. Colors encode the nature of the underlying magnetic state: green and yellow for the compressible Y-like and V-like supersolid phases, and blue and light blue for the incompressible UUD phase corresponding to a $m=1/3$ plateau. Black lines are linear fits to the data using the averaged $g_{z}$ values, see SI Sec. IC for details. Dotted lines represent excitations originating from the Y and V phases, solid lines indicate dominant spin-flip modes in the UUD phase, and dashed lines correspond to disorder-induced satellite modes of the UUD phase.
@@
At first glance, the magnetization process of antiferromagnets is simple. In the paramagnetic regime, Langevin or Brillouin theory describes the response of independent spins $S$—treated either as classical dipoles or quantum moments—to temperature $T$ and applied magnetic field $H$ [12]. In the magnetically ordered state, assuming that spins initially lie in the plane perpendicular to the magnetic field direction $z$, the classical dipoles cant toward the field direction. This induces a uniform magnetization $M(H)$ which increases linearly until $\langle S^{z}\rangle=S$ for each spin (Fig. 1a). The onset of this fully polarized state defines the saturation field $H_{\rm sat}$. A closer inspection reveals that non-linear contributions to the isothermal magnetization encode rich phenomena. For example, quantum corrections result in a convex magnetization in Heisenberg $S\!=\!1/2$ antiferromagnets [13, 14]. In geometrically frustrated systems, fractional magnetization plateaus can occur [15, 16, 17] with $\langle\sum_{i}S_{i}^{z}\rangle=\langle S^{z}_{\rm tot}\rangle\!=\!m\!\times\!NS$, where $m$ is a rational number and $N$ is the total number of spins (Fig. 1b). In systems with biquadratic exchange, quantum and classical effects compete in the stabilization of magnetization plateaus [18] while in frustrated $S\!>\!1/2$ systems with uniaxial anisotropy, a Devil’s staircase with an infinite number of magnetization plateaus is possible at the classical level [19]. Clearly, magnetization plateaus are a widespread and fascinating occurrence in magnetic systems [20]. They reflect that the total spin operator is conserved and commutes with the Hamiltonian, i.e., $[S^{z}_{\rm tot},\mathcal{H}]\!=\!0$. As the resulting magnetic excitations are gapped, a reliable determination of the system Hamiltonian $\mathcal{H}$ [21, 22] is possible at much lower magnetic fields than those necessary to reach a fully polarized state. Furthermore, regardless of the underlying classical or quantum mechanisms that stabilize the plateau, the resulting magnetic state is incompressible [20]; an increase in $H$ does not lead to an increase in $M(H)$. This incompressibility creates a highly sensitive platform for detecting even minute amounts of chemical disorder, because spins near a defect can respond strongly to the magnetic field, i.e. these spins are compressible.
@@
To demonstrate this spectacular effect, we focus on the Heisenberg-Ising (or XXZ) triangular-lattice antiferromagnet K2Co(SeO3)2 (KCSO). This compound has long been known [23, 24] and was recently revisited in detail [25]. It features a robust magnetization plateau that spans an extraordinarily wide field range, from $2$ to $17$ T [26]. The magnetization plateau is flanked by compressible spin supersolid phases at low temperature [26, 27, 28]. These phases are the quantum analogs of non-collinear Y and V phases, where a static magnetic order in the $z$-direction coexists with phase coherence—but no static order—transverse to $z$. Although the exact nature of these phases is not yet fully understood, we will refer to them as Y and V in the rest of this article. The crystal structure of KCSO, depicted in Fig. 1c, comprises isolated $\rm CoO_{6}$ octahedra connected by $\rm SeO_{3}$ tetrahedra, forming a triangular lattice of $\rm Co^{2+}$ ions in the $ab$-plane. These triangular layers are stacked along the $c$-axis and separated by layers of K ions. Although the effective magnetic moments of KCSO extracted from the Curie-Weiss fit at high temperatures are nearly isotropic, with $\mu_{\rm eff}^{c}=6.14\mu_{\rm B}\approx 1.13\mu_{\rm eff}^{ab}$ (where $\mu_{\rm B}$ is the Bohr magneton) corresponding to an effective $g$-factor of $g_{c}=7.1$ [25], bilinear interactions exhibit a strong Ising character. The out-of-plane nearest-neighbor exchange $J_{zz}$ is an order of magnitude larger than the in-plane exchange $J_{xy}$ [26, 27]. Above the critical field $\mu_{0}H_{\rm c1}\approx 2$ T (where $\mu_{0}$ is the vacuum permeability) and at or below $T=5$ K, KCSO stabilizes an archetypal up-up-down (UUD) phase associated with a magnetization plateau at $m\!=\!1/3$. This state persists until the critical field $\mu_{0}H_{\rm c2}\approx 17$ T.
@@
Figure 2: Schematic interpretation of magnetic excitations in the UUD phase of K2Co(SeO3)2. a, Magnetic excitations expected in the UUD phase with and without bond disorder. Solid lines (labeled #1 and #2) represent excitations in a clean system, while dashed lines (labeled #1a for above, and #1b, #2b for below) correspond to additional excitations in a disordered system. Blue and light blue indicate major spin-flip processes, while red and orange represent minor spin-flip processes. b, Spin structure of the ideal UUD phase for a $3\times 3$ triangular-lattice supercell. c, Visualization of major and minor spin-flip processes in clean and disordered systems. The flipped spin in each process is shown with a green sphere. The green bond represents a disorder bond with modified exchange interactions $J_{zz}\textquoteright=(1-d_{zz})J_{zz}$ and $J_{xy}\textquoteright=(1-d_{xy})J_{xy}$.
@@
Here, we utilize far-infrared magneto-optical spectroscopy (FIRMS) and thermomagnetic measurements to investigate the behavior of KCSO near and within the plateau at $m=1/3$ (see Methods). To our surprise, although our crystals are grown using published protocols and chemical disorder ruled out by advanced X-ray diffraction and electron microscopy methods (see SI Secs. IV and V), the measurements reveal sharp spectroscopic fingerprints that can be clearly linked to very low levels of chemical disorder. These fingerprints were not seen in previous measurements [27]. Our optical measurements in the UUD phase reveal two dominant modes accompanied by three satellite modes, as illustrated in Fig. 1d and 1e. These satellite modes cannot be explained by clean-system models across three simulation techniques: linear spin wave theory (LSWT), Landau-Lifshitz dynamics (LLD), and exact diagonalization (ED). Instead, we develop a disordered-system model that fully captures our experimental observations. In this framework, the energy shift between the dominant and satellite modes is proportional to the strength of exchange disorder on nearest-neighbor bonds. The large energy shifts in our data suggest that the disorder in our KCSO crystals is more akin to spin vacancies than bond disorder. A typically undetectable amount of vacancies—as little as 2.5%—is sufficient to produce the observed intensity of the satellite modes. Moreover, magnetization and heat capacity measurements reveal a secondary magnetization plateau state with $m\!<\!1/3$, which is successfully reproduced in classical Monte Carlo (MC) simulations using the vacancy model.
@@
Our FIRMS results, measured in Faraday geometry ($H\parallel c\equiv z$) at $T\!=\!5$ K, are shown in Fig. 1d. At low magnetic fields ($30<\mu_{0}H<$ 3 T), the system is in the Y phase [27], and the spectra exhibit two broad field-dependent modes that cross at $\mu_{0}H\approx$ 3 T. Around 3 T, the system enters the UUD phase and displays much sharper modes, with the two strongest intersecting at $\mu_{0}H\!=\!9.75$ T. Three satellite modes, running parallel to the main modes, are also observed. Above 15 T, an additional mode emerges and broadens with increasing field; we associate this mode with the onset of the V phase [27]. We determine the energy of major modes at each magnetic field, $E(H)$, using Lorentzian fits, and plot the results on Fig. 1e, where the color encodes the parent Y, UUD, or V phase associated with each mode. Now, we address the origin of the two strongest modes in the UUD phase. The magnetic Hamiltonian of KCSO in that field configuration reads
@@
$E_{n}=nJ_{zz}\pm\mu_{0}\mu_{\rm B}g_{z}H,~{}~{}(n=0,1,2,3,…).$ (2)
@@
Due to optical selection rules, only transitions with $\Delta S^{z}=\pm 1$ are visible in FIRMS. Therefore, our experiment primarily detects excited states involving single or triple spin-flip processes (see SI Sec. II for further discussion). The two strongest modes observed in the UUD phase can be naturally attributed to single spin-flip excitations. In Fig. 2a, solid lines represent the mode energies $E(H)$ expected for single spin-flips allowed by the selection rule. These processes can occur at two distinct spin sites: up spin (major spin) and down spin (minor spin). A minor spin flip costs onsite energy in terms of $J_{zz}$ only—specifically, $3J_{zz}$ for the triangular lattice—since the flipped spin cannot propagate to neighboring spin sites. In contrast, a major spin flip incurs no onsite energy change in the classical limit but is accompanied by an additional quantum correction of $23J_{xy}/$ 2, reflecting that the flipped spin can propagate to three neighboring up-spin sites (Fig. 2b,c). Linear fits to these two modes (see SI Tab. S2) yield $J_{xy}$ = 0.20(1) meV, $J_{zz}$ = 2.88(1) meV, and $g_{z}$ = 7.43 (positive slope) or $g_{z}$ = 7.50 (negative slope), in good agreement with previous neutron scattering and magnetization studies [25, 26, 27] but with a more accurate determination of the $g$-factor. Triple spin-flip excitations, forming both bound-state and a free particle continuum, are also observed (see SI Sec . IIA-B for details).
$J_{zz}^{\prime}$ $=$ $J_{zz}(1-d_{zz})~{}~{}{\rm and}$ (3)
$J_{xy}^{\prime}$ $=$ $J_{xy}(1-d_{xy}),$ (4)
where $d_{zz}$ and $d_{xy}$ represent the disorder-induced reduction of the respective interaction. Notably, this model yields fine spectroscopic features that are the unmistakable hallmark of the disorder’s nature. Specifically, a major spin-flip excitation can split into two modes, depending on whether the disorder bond is connected to a major spin (mode #1a in Fig. 2) or a minor spin (mode #1b in Fig. 2). The excitation energies are shifted by ±$\pm J_{zz}d_{zz}/2$, respectively. Moreover, when the disorder bond is connected to a major spin, there is a quantum correction of $-\alpha J_{xy}d_{xy}$ (where $\alpha\simeq n_{dm}$ is the disorder density in the dilute limit), due to the propagation of the flipped spin to two neighboring up-spin sites with intact bonds. In the case of a minor spin-flip process, the disorder only introduces an energy shift of $-J_{zz}d_{zz}/2$, since the flipped spin cannot propagate to neighboring spin sites (see detailed discussion in SI Sec. IID). More generally, we hypothesize that the splitting and multiplicity of spin-flip excitations in incompressible magnetic states serve as a unique indicator of bond disorder.
@@
Figure 4: Magnetization and specific heat of our crystals of K2Co(SeO3)2. a, Temperature dependence of the magnetization of KCSO in a $c$-axis magnetic field. b-c, Simulated isothermal magnetization curves using the classical MC method for clean and disordered systems, respectively. The color of each curve corresponds to the temperature shown on the color bar. d, Schematic mechanism stabilizing an additional magnetization plateau around $\mu_{0}H=2.8$ T. The top and bottom panels show the spin configurations at 2 T and 4 T, respectively. Red, blue, and white circles indicate up spins, down spins, and vacancies, respectively. Orange circles highlight spin-flip sites that result in the additional plateau. e, Field derivative of the magnetization. f-g, Temperature and magnetic field dependence of the specific heat divided by temperature for KCSO in a $c$-axis magnetic field. The colorplot combines constant-field ($T\geq 0.8$ K) and constant-temperature ($T<0.8$ K) measurements.
@@
Using the analytical solution of the disorder-bond model, we perform linear fits to the observed satellite modes (see SI Sec. IC). We find that the $y$-intercepts of these three modes are consistent with the model prediction, as illustrated in Fig. 2a. Quantitatively, we obtain $d_{zz}\!=\!0.93$ from the energy shift between modes #1 and #1b, and $d_{zz}=0.96$ from the energy shift between modes #2 and #2b. The close-to-unity value for $d_{zz}$ points to the complete absence of a bond, rather than a mere modification of the exchange due to disorder. The simplest mechanism to generate missing bonds is through spin-site vacancies, i.e., missing $\rm Co^{2+}$ ions.

### v3

# Sharp spectroscopic fingerprints of disorder in an incompressible magnetic state

@@
Disorder significantly impacts the electronic properties of conducting quantum materials by inducing electron localization and thus altering the local density of states and electric transport. In insulating quantum magnetic materials, the effects of disorder are less understood and can drastically impact fluctuating spin states like quantum spin liquids. In the absence of transport tools, disorder is typically characterized using chemical methods or by semi-classical modeling of spin dynamics. This requires high magnetic fields that may not always be accessible. Here, we show that magnetization plateaus—incompressible states found in many quantum magnets—provide an exquisite platform to uncover small amounts of disorder, regardless of the origin of the plateau. Using optical magneto-spectroscopy on the Ising-Heisenberg triangular-lattice antiferromagnet K2Co(SeO3)2 exhibiting a 1/3 magnetization plateau, we identify sharp spectroscopic lines, the fine structure of which serves as a hallmark signature of disorder. Through analytical and numerical modeling, we show that these fingerprints not only enable us to quantify minute amounts of disorder but also reveal its nature—as dilute vacancies. Remarkably, this model explains all details of the thermomagnetic response of our system, including the existence of multiple plateaus. Our findings provide a new approach to identifying disorder in quantum magnets.
Introduction
@@
The role of disorder in quantum magnetic materials is less understood yet it is crucial for a wide range of systems, particularly quantum spin liquids (QSLs). These systems are believed to host topological order and itinerant fractional excitations without exhibiting clear, conventional order-parameter-like signatures. The response, survival, and nature of potential QSLs in the presence of chemical disorder are not fully understood. Increasingly, researchers are recognizing that disorder can mimic the expected behavior of a QSL [43, 15, 16, 14, 28, 19]. But a significant challenge lies in quantifying the level of disorder using chemical techniques and translating this information into a model magnetic Hamiltonian, $\mathcal{H}$, whose emergent quantum many-body phases are not known in advance. A successful approach involves employing electron spin resonance [17, 22] or fitting inelastic neutron scattering and optical spectroscopy measurements under sufficiently high magnetic fields to polarize the system. In this semi-classical regime, it becomes possible to directly measure $\mathcal{H}$ [7], including the level of chemical disorder [20, 27]. However, practical limitations arise from the need for large magnetic fields. Furthermore, spin-anisotropic systems of current interest may only reach saturation asymptotically [41].
@@
The complex magnetization process of antiferromagnets may offer an alternative strategy. By utilizing quantized magnetization plateaus—when they occur—we show that it is not only possible to extract $\mathcal{H}$ at a much lower magnetic field, but these plateaus also exhibit unique, sharp spectroscopic signatures allowing for the detection of disorder with high precision. Similar to the QHE, this sensitivity emerges from the distinctive structure of incompressible magnetic states with quantized magnetization, an effect which has not been widely acknowledged until now.
@@
Figure 1: Magnetization plateau and far-infrared magneto-spectroscopy measurements of K2Co(SeO3)2. a, Schematic representation of the normalized uniform magnetization $m=\langle S_{\rm tot}^{z}\rangle/NS$ of an ordered antiferromagnet described at the classical level (solid line) or including quantum corrections (dashed line). The spin-$S$ moments form a canted structure and become fully polarized at the saturation magnetic field $H_{\rm sat}$. b, Normalized uniform magnetization of an antiferromagnet with a field-induced quantized plateau state, which is stabilized between $H_{\rm c1}$ and $H_{\rm c2}$. The UUD state is a possible classical state realizing a $m=1/3$ magnetization plateau. c, Crystal structure of KCSO. Orange spheres represent oxygen, purple represents potassium, blue represents cobalt, and yellow represents selenium ions. The triangular lattice of magnetic Co2+ ions is shown by blue bonds. d, Frequency and magnetic field dependence of the normalized magneto-transmission spectra $I(\omega;H)$ of KCSO measured in Faraday geometry at T=5T\!=\!5 K with the magnetic field direction $z$ along the $c$-axis, see Methods and Supplementary Sec. IA-B for details of our measurement and data analysis protocols. e, Energy modes $E(H)$ extracted from Lorentzian fits to the spectra at each magnetic field. Colors encode the nature of the underlying magnetic state: green and yellow for the compressible Y-like and V-like supersolid phases, and blue, light blue, and purple for the incompressible UUD phase corresponding to a $m=1/3$ plateau. Black lines are linear fits to the data using the averaged $g_{z}$ values, see Supplementary Sec. IC for details. Dotted lines represent excitations originating from the Y and V phases, solid lines indicate dominant spin-flip modes, dot-dashed lines show the triple spin-flip modes, and dashed lines correspond to disorder-induced satellite modes of the UUD phase.
@@
At first glance, the magnetization process of antiferromagnets is simple. In the paramagnetic regime, Langevin or Brillouin theory describes the response of independent spins SS—treated either as classical dipoles or quantum moments—to temperature $T$ and applied magnetic field $H$ [3]. In the magnetically ordered state, assuming that spins initially lie in the plane perpendicular to the magnetic field direction $z$, the classical dipoles cannot move toward the field direction. This induces a uniform magnetization $M(H)$ which increases linearly until $\langle S^{z}\rangle=S$ for each spin (Fig. 1a). The onset of this fully polarized state defines the saturation field $H_{\rm sat}$. A closer inspection reveals that non-linear contributions to the isothermal magnetization encode rich phenomena. For example, quantum corrections result in a convex magnetization in Heisenberg S=1/2S\!=\!1/2 antiferromagnets [4, 39]. In geometrically frustrated systems, fractional magnetization plateaus can occur [6, 38, 21] with $\langle\sum_{i}S_{i}^{z}\rangle=\langle S^{z}_{\rm tot}\rangle\!=\!m\!\times\!NS$, where $m$ is a rational number and $N$ is the total number of spins (Fig. 1b). In systems with biquadratic exchange, quantum and classical effects compete in the stabilization of magnetization plateaus [1] while in frustrated S>1/2S\!>\!1/2 systems with uniaxial anisotropy, a Devil’s staircase with an infinite number of magnetization plateaus is possible at the classical level [2]. Clearly, magnetization plateaus are a widespread and fascinating occurrence in magnetic systems [29]. They reflect that the total spin operator is conserved and commutes with the Hamiltonian, i.e., $S^{z}_{\rm tot},\mathcal{H}]\!=\!0$. As the resulting magnetic excitations are gapped, a reliable determination of the system Hamiltonian $\mathcal{H}$ [13, 36] is possible at much lower magnetic fields than those necessary to reach a fully polarized state. Furthermore, regardless of the underlying classical or quantum mechanisms that stabilize the plateau, the resulting magnetic state is incompressible [29]; an increase in $H$ does not lead to an increase in $M(H)$. This incompressibility creates a highly sensitive platform for detecting even small amounts of chemical disorder, because spins near a defect can respond strongly to the magnetic field, i.e., these spins are compressible.
@@
To demonstrate this spectacular effect, we focus on the Ising-Heisenberg (or XXZ) triangular-lattice antiferromagnet K2Co(SeO3)2 (KCSO). This compound has long been known [35, 34] and was recently revisited in detail [40]. It features a robust magnetization plateau that spans an extraordinarily wide field range, from $2$ to $17$ T [42]. The magnetization plateau is flanked by compressible spin supersolid phases at low temperature [42, 5, 37]. These phases are the quantum analogs of non-collinear Y and V phases, where a static magnetic order in the $z$-direction coexists with phase coherence—but no static order—transverse to $z$. Although the exact nature of these phases is not yet fully understood, we will refer to them as Y and V in the rest of this article. The crystal structure of KCSO, depicted in Fig. 1c, comprises isolated $\rm CoO_{6}$ octahedra connected by $\rm SeO_{3}$ tetrahedra, forming a triangular lattice of $\rm Co^{2+}$ ions in the $ab$-plane. These triangular layers are stacked along the $c$-axis and separated by layers of K ions. Although the effective magnetic moments of KCSO extracted from the Curie-Weiss fit at high temperatures are nearly isotropic, with $\mu_{\rm eff}^{c}=6.14\mu_{\rm B}\approx 1.13\mu_{\rm eff}^{ab}$ (where $\mu_{\rm B}$ is the Bohr magneton) corresponding to an effective $g$-factor of $g_{c}=7.1$ [40], bilinear interactions exhibit a strong Ising character. The out-of-plane nearest-neighbor exchange $J_{zz}$ is an order of magnitude larger than the in-plane exchange $J_{xy}$ [42, 5]. Above the critical field $\mu_{0}H_{\rm c1}\approx 2$ T (where $\mu_{0}$ is the vacuum permeability) and at or below $T=5$ K, KCSO stabilizes an archetypal up-up-down (UUD) phase associated with a magnetization plateau at m=1/3m\!=\!1/3. This state persists until the critical field $\mu_{0}H_{\rm c2}\approx 17$ T.
@@
Figure 2: Schematic interpretation of magnetic excitations in the UUD phase of K2Co(SeO3)2. a, Magnetic excitations expected in the UUD phase with and without bond disorder. Solid lines (labeled #1 and #2) represent excitations in a clean system, while dashed lines (labeled #1a for above, and #1b, #2b for below) correspond to additional excitations in a disordered system. Blue and light blue indicate major spin-flip processes, while red and orange represent minor spin-flip processes. b, Spin structure of the ideal UUD phase for a $3\times 3$ triangular-lattice supercell. c, Visualization of major and minor spin-flip processes in clean and disordered systems. The flipped spin in each process is shown with a green sphere. The green bond represents a disorder bond with modified exchange interactions $J_{zz}’=J_{zz}(1-d_{zz})$ and $J_{xy}’=J_{xy}(1-d_{xy})$.
Results
@@
Here, we utilize far-infrared magneto-optical spectroscopy (FIRMS) and thermomagnetic measurements to investigate the behavior of KCSO near and within the plateau at $m=1/3$ (see Methods). To our surprise, although our crystals are grown using published protocols and chemical disorder ruled out by advanced X-ray diffraction and electron microscopy methods (see Supplementary Sec. V and VI), the measurements reveal sharp spectroscopic fingerprints that can be clearly linked to very low levels of chemical disorder. These fingerprints were not seen in previous measurements [5]. Our optical measurements in the UUD phase reveal two dominant modes accompanied by three satellite modes, as illustrated in Fig. 1d and 1e. These satellite modes cannot be explained by clean-system models across three simulation techniques: linear spin wave theory (LSWT), Landau-Lifshitz dynamics (LLD), and exact diagonalization (ED). Instead, we develop a disordered-system model that fully captures our experimental observations. In this framework, the energy shift between the dominant and satellite modes reflects the nature and strength of exchange disorder on nearest-neighbor bonds. The large and discrete energy shifts in our data suggest that the disorder in our crystals is akin to missing bonds or spin vacancies, rather than a distribution of weakened bonds. As we illustrate below, a typically small amount of vacancies—as little as 2.3%—is sufficient to produce the observed intensity of the satellite modes. Moreover, magnetization and heat capacity measurements reveal a secondary magnetization plateau state with m<1/3m\!<\!1/3, which is successfully reproduced in classical Monte Carlo (MC) simulations using the vacancy model.
@@
Our FIRMS results, measured in Faraday geometry ($H\parallel c\equiv z$) at T=5T\!=\!5 K, are shown in Fig. 1d. At low magnetic fields ($0<\mu_{0}H<3$ T), the system is in the Y phase [5], and the spectra exhibit two broad field-dependent modes that cross at $\mu_{0}H\approx$ 3 T. Around 3 T, the system enters the UUD phase and displays much sharper modes, with the two strongest intersecting at $\mu_{0}H\!=\!9.75$ T. Three satellite modes, running parallel to the main modes, are also observed. Above 15 T, an additional mode emerges and broadens with increasing field; we associate this mode with the onset of the V phase [5]. We determine the energy of major modes at each magnetic field, $E(H)$, using Lorentzian fits, and plot the results on Fig. 1e, where the color encodes the parent Y, UUD, or V phase associated with each mode. Now, we address the origin of the two strongest modes in the UUD phase. The magnetic Hamiltonian of KCSO in that field configuration reads
@@
[…]
Due to optical selection rules, only transitions with $\Delta S^{z}=\pm 1$ are visible in FIRMS. Therefore, our experiment primarily detects excited states involving single or triple spin-flip processes (see Supplementary Sec. II for further discussion). The two strongest modes observed in the UUD phase can be naturally attributed to single spin-flip excitations. In Fig. 2a, solid lines represent the mode energies $E(H)$ expected for single spin-flips allowed by the selection rule. These processes can occur at two distinct spin sites: up spin (major spin) and down spin (minor spin). A minor spin flip costs onsite energy in terms of $J_{zz}$ only—specifically, $3J_{zz}$ for the triangular lattice—since the flipped spin cannot propagate to neighboring spin sites. In contrast, a major spin flip incurs no onsite energy change in the classical limit but is accompanied by an additional quantum correction of $3J_{xy}/2$, reflecting that the flipped spin can propagate to three neighboring up-spin sites (Fig. 2b,c). Linear fits to these two modes (see Supplementary Table 2) yield $J_{xy}$ = 0.20(1) meV, $J_{zz}$ = 2.88(1) meV, and $g_{z}$ = 7.45(5), in good agreement with previous neutron scattering and magnetization studies [40, 42, 5], but providing a more accurate determination of the $g$-factor. Remarkably, in the UUD phase, all main and satellite modes appear resolution-limited, as shown in Supplementary Figure 4. Triple spin-flip excitations, forming both a bound-state and a free particle continuum, are also observed (see Supplementary Sec. IIA for details).
$J_{zz}^{\prime}=J_{zz}(1-d_{zz}),\quad J_{xy}^{\prime}=J_{xy}(1-d_{xy}),$ (3)
where $d_{zz}$ and $d_{xy}$ represent the disorder-induced reduction of the respective interaction. Notably, this model yields fine spectroscopic features that are the unmistakable hallmark of the disorder’s nature. Specifically, a major spin-flip excitation can split into two modes, depending on whether the disorder bond is connected to a major spin (mode #1a in Fig. 2) or a minor spin (mode #1b in Fig. 2). The excitation energies are shifted by $\pm J_{zz}d_{zz}/2$, respectively. Moreover, when the disorder bond is connected to a major spin, there is a quantum correction of $\alpha J_{xy}d_{xy}$ (where $\alpha\simeq n_{dm}$ is the disorder density in the dilute limit), due to the propagation of the flipped spin to two neighboring up-spin sites with intact bonds. In the case of a minor spin-flip process, the disorder only introduces an energy shift of $J_{zz}d_{zz}/2$, since the flipped spin cannot propagate to neighboring spin sites (see detailed discussion in Supplementary Sec. IID). More generally, we hypothesize that the splitting and multiplicity of spin-flip excitations in incompressible magnetic states serve as a unique indicator of bond disorder.
Figure 3: Universality of the disorder-bond model and its correspondence with the vacancy model. Schematic UUD spin configurations and excitation modes for: a-b, the weakened-bond model, where for illustration the exchange interactions on green and dark green bonds are reduced with parameters $d_{zz}=d_{xy}=0.1$ (green) and $d_{zz}=d_{xy}=0.2$ (dark green); c-d, the broken-bond model, where $d_{zz}=d_{xy}\rightarrow 1$ for the same broken bond; and e-f, the vacancy model, where two magnetic sites are removed. Throughout, red (blue) circles indicate up (down) spins. The line style for the spectra in b, d, and f follows the convention of Fig. 2. The spectral intensity of the satellite modes is not shown and differs between panels d and f, which represent different defect densities.
@@
Figure 5: Magnetization and specific heat of our crystals of K2Co(SeO3)2. a, Temperature dependence of the magnetization of KCSO in a $c$-axis magnetic field. b-c, Simulated isothermal magnetization curves using the classical MC method for clean and disordered systems, respectively. Magenta dot-dashed lines in each figure indicate the magnetization of the pre-plateau and UUD plateau phase with the vacancy concentration $n$. d, Schematic mechanism stabilizing an additional magnetization plateau around $\mu_{0}H=2.8$ T. The top and bottom panels show the spin configurations at 2 T and 4 T, respectively. Red, blue, and white circles indicate up spins, down spins, and vacancies, respectively. Orange circles highlight spin-flip sites that result in the additional plateau. e, Field derivative of the magnetization. f-g, Temperature and magnetic field dependence of the specific heat divided by temperature for KCSO in a $c$-axis magnetic field. The colorplot combines constant-field ($T\geq 0.8$ K) and constant-temperature ($T<0.8$ K) measurements.
@@
Using the analytical solution of the disorder-bond model, we perform linear fits to the observed satellite modes (see Supplementary Sec. IC). We find that the $y$-intercepts of these three modes are consistent with the model prediction, as illustrated in Fig. 2a. Quantitatively, we obtain $d_{zz}\!=\!0.93$ from the energy shift between modes #1 and #1b, and $d_{zz}=0.96$ from the energy shift between modes #2 and #2b. The close-to-unity value for $d_{zz}$ points to the complete absence of a bond, rather than a mere modification of the exchange due to disorder. The simplest mechanism to generate missing bonds is through spin-site vacancies, i.e., missing $\rm Co^{2+}$ ions.
The disorder-bond framework of Eq. 3 interpolates between three distinct models of random disorder: a weakened-bond model, a broken-bond model, and a vacancy model. Figure 3 schematically illustrates the spin and exchange configurations for each case and the associated excitation spectra. In the weakened-bond model, multiple bonds exhibit randomly distributed reduction factors $d_{zz}$ and $d_{xy}$. The resulting satellite excitations detach from the main modes, with their energy offsets set by the reduction factors. Consequently, for a distribution of weakened bonds, we expect the main spectral lines to be broadened and sharp satellite modes to be absent, depending on the details of the disorder distribution (see Fig. 3b). By contrast, in the broken-bond model, the reduction factors are fixed to unity on each defective bond. This produces sharp satellite modes well separated from the main modes, which also remain sharp (see Fig. 3d). Remarkably, the vacancy model yields an identical spectrum, albeit with distinct intensity, because the six bonds adjacent to a vacancy are effectively removed (see Fig. 3e). Thus, despite their distinct chemical origins, the broken-bond and vacancy models are spectroscopically indistinguishable for an unknown defect density. From a materials standpoint, however, the complete absence of a specific bond is less realistic than the presence of a magnetic-ion vacancy.